### v1

# MiniScope: Automated UI Exploration and Privacy Inconsistency Detection of MiniApps via Two-phase Iterative Hybrid Analysis

Abstract.
The advent of MiniApps, operating within larger SuperApps, has revolutionized user experiences by offering a wide range of services without the need for individual app downloads. However, this convenience has raised significant privacy concerns, as these MiniApps often require access to sensitive data, potentially leading to privacy violations. Our research addresses the critical gaps in the analysis of MiniApps’ privacy practices, especially focusing on WeChat MiniApps in the Android ecosystem. Despite existing privacy regulations and platform guidelines, there is a lack of effective mechanisms to safeguard user privacy fully.
We introduce MiniScope, a novel two-phase hybrid analysis approach, specifically designed for the MiniApp environment. This approach overcomes the limitations of existing static analysis techniques by incorporating dynamic UI exploration for complete code coverage and accurate privacy practice identification. Our methodology includes modeling UI transition states, resolving cross-package callback control flows, and automated iterative UI exploration. This allows for a comprehensive understanding of MiniApps’ privacy practices, addressing the unique challenges of sub-package loading and event-driven callbacks. Our empirical evaluation of over 120K MiniApps using MiniScope demonstrates its effectiveness in identifying privacy inconsistencies. The results reveal significant issues, with 5.7% of MiniApps over-collecting private data and 33.4% overclaiming data collection. These findings emphasize the urgent need for more precise privacy monitoring systems and highlight the responsibility of SuperApp operators to enforce stricter privacy measures.
MiniApps, Privacy Compliance
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Software security engineering
1. Introduction
Mini-programs or MiniApps, epitomizing a novel breed of lightweight mobile applications, have been transformative in the realm of user experience in recent years (Yang et al., 2022b). MiniApps are lightweight versions of full-fledged applications which operate within a host application or “SuperApp”. They can provide users with a diverse range of services under the “SuperApp + MiniApps” business paradigm umbrella (Kumar et al., 2018). From online shopping, gaming, and accessing social media to availing healthcare services (wec, 2022; Rao and Ko, 2021; Hao et al., 2018; Qian and Hanser, 2021), these MiniApps offer the richness of standalone applications without necessitating individual downloads. This seamless integration has radically enhanced user experiences, with simplified navigation and easy-to-access features all within a single app interface. Currently, the most representative SuperApp, WeChat, hosts a staggering 3.5 million MiniApps that engage over 600 million daily active users (Yang et al., 2023). Its extensive ecosystem and unprecedented user engagement underscore the revolutionary impact of the MiniApp paradigm, which could introduce complications and unexpected privacy concerns at the same time.
To deliver their diverse services, these MiniApps often require access to sensitive system resources (e.g., camera and Bluetooth), as well as user data (e.g., location information, phone numbers, and email accounts). This has given rise to potential privacy violations, as third-party developers within the SuperApp ecosystem could access and utilize this information, often without clear or explicit user consent. Accompanied by the rising awareness and concern for user privacy, numerous regulations and laws, including the General Data Protection Regulation (GDPR) (GDP, 2022), California Consumer Privacy Act (CCPA) (CCP, 2022), Act on the Protection of Personal Information (APPI) (APP, 2022), and Canadian Consumer Privacy Protection Act (CPPA) (CPP, 2022), have been enforced to ensure the transparent and accountable collection and processing of user data. Meanwhile, SuperApp platforms, including WeChat, have also incorporated guidelines and instructions that demand stricter scrutiny of data collection and processing from MiniApps (Wec, 2023d). However, these measures, while laudable, have proven insufficient to fully safeguard the privacy of users within the expansive SuperApp ecosystems (Wec, 2023a, b). Considering the openness of Android ecosystem and numerous prior studies focuses on WeChat MiniApps in Android (Wang et al., 2023b, a; Zhang et al., 2023a), our work particularly focuses on the same target, i.e., the WeChat MiniApp in Android, unless stated otherwise.
Research Gaps. While considerable progress have been made in the field of privacy inconsistency analysis for mobile apps (Slavin et al., 2016; Wang et al., 2018; Zimmeck et al., 2016) and Web apps (Ling et al., 2022; Bui et al., 2023), the unique features of MiniApps present distinct challenges that remain largely unexplored. Existing techniques, primarily centered around taint analysis (Arzt et al., 2014; Enck et al., 2014; Wang et al., 2023b; Meng et al., 2023; Li et al., 2023), encounter two major limitations when adapted to the MiniApp environment. Firstly, unlike traditional app packaging mechanisms that allows access to all source code at once, MiniApps employ a subpackaging mechanism (W3C, 2023b), that is, only the main package is initialized during a cold start and the sub-packages are loaded on demand during runtime. Consequently, methods that rely solely on static analysis (Wang et al., 2023b; Meng et al., 2023; Li et al., 2023) fail to trigger the loading of these sub-packages, thereby overlooking a substantial amount of code that could involve privacy practices. This oversight leads to an incomplete data flow graph, resulting in a higher incidence of false negatives in taint analysis. Secondly, MiniApps are framework-based and event-driven. Prior research (Yang et al., 2015; Liu et al., 2022, 2023) has demonstrated the critical importance of user-driven callback control flow analysis in similar application paradigm. However, existing taint analysis methods (Wang et al., 2023b; Meng et al., 2023; Li et al., 2023) concentrate primarily on sensitive data flows while neglecting event-driven callbacks, falling short in thoroughly scrutinizing asynchronous entry points of MiniApp logics. This neglect results in a significant number of false positives in taint analysis, as some taint paths obtained could be potential unreachable.
Technical Challenges. Given the above research gaps, the challenges we faced are twofold. The static analysis of MiniApps presents difficulties due to the need for sound resolution of user interactions and event-driven callbacks. Although similar techniques have already been widely explored in Android/iOS (Yang et al., 2015; Liu et al., 2022, 2023), the unique challenge in constructing control flow in MiniApps lies in the cross-file and cross-subpackage callbacks resolution. Moreover, the static analysis is intricately intertwined with dynamic UI exploration. That is, in the context of MiniApps, the prerequisite for sound static analysis is to load sub-package on demand through UI exploration, while efficient UI exploration requires static prior knowledge (such as UI transition states) for guidance. Therefore, this presents a logical paradox, which demands designing tailored hybrid analysis strategies specifically for MiniApps.
Our Approach. In response to the research gaps and challenges in analyzing MiniApps, we introduce MiniScope, an innovative two-phase iterative hybrid analysis approach tailored for in-depth UI exploration and precise privacy practices identification. The initial phase of our approach focuses on dynamic sub-package loading through UI exploration to obtain the full source code. Utilizing a static analyzer, we initialize a MiniApp Dependency Graph (MDG) for the main package, meticulously designed to extract UI transition states, event-driven control flows, and data flows. This initial MDG serves as the groundwork for accurate taint analysis and guides the subsequent UI exploration. Based on the initial MDG, the directed UI explorer then engages in fuzzy matching between WXML components and the UI Widgets Tree. By adopting a sub-package-directed breadth-first traversal strategy, MiniScope dynamically explore sub-package pages, thereby obtaining full source code of the MiniApp. In the second phase, our focus shifts to the precise identification of privacy practices. MiniScope merges the main package with the dynamically loaded sub-packages, performing static analysis to obtain the complete MDG. In this phase, the directed UI explorer utilizes a privacy-practice-directed depth-first strategy for runtime exploration of sensitive behaviors. MiniScope then cross-verifies the privacy practices extracted from both static and dynamic analysis against the declared privacy policy, enabling a thorough detection to identify any discrepancies or privacy inconsistencies at a granular level.
To demonstrate the effectiveness practically, we utilize MiniScope to detect privacy inconsistency and carry out a comprehensive evaluation involving over 120K MiniApps. Following a filtration process to exclude those without a valid privacy policy, we eventually analyze 10,786 MiniApps. The results underscore the superior performance of MiniScope in identifying privacy-related practices when compared with TaintMini (Wang et al., 2023b), the state-of-the-art technique. Specifically, MiniScope demonstrate 7.9% and 23.5% improvement in precision and recall respectively. Our analysis further reveal that 5.7% of the MiniApps are secretly over-collecting private data, and 33.4% of the MiniApps overclaim the data they actually collect. We have responsibly disclosed our findings to 2,282 developers, receiving 44 confirmations and acknowledgments. These findings highlight the pressing need for the implementation of a more precise privacy monitoring system, emphasizing the responsibility of SuperApp operators to enforce such measures.
Contributions. Our contributions are summarized as follows:
[…]
2. Background
2.1. Features of MiniApps
Architecture of MiniApps. The runtime environment of a MiniApp (Figure 1), facilitated by SuperApps like WeChat or Alipay, is characterized by a dual-thread architecture splitting into a render (or view) layer and a logic layer (Wang et al., 2023b). The render layer, akin to HTML and CSS, involves WXML (markup language) and WXSS (style sheet), while the logic layer consists of JavaScript files. MiniApp pages correspond to multiple WebView threads in the render layer, whereas logical operations, data requests, and interface calls occur via JSCore threads in the logic layer. The two layers engage through WeChat’s JSBridge mechanism, handling user interaction events and data updates. Furthermore, the logic layer introduces sub-app APIs like wx.getLocation for user location retrieval, which are bridged into the SuperApp’s native layer (Lu et al., 2020).
Figure 1. Architecture of MiniApps.
Sub-package Dynamic Loading. MiniApps employ a unique sub-package mechanism to stay lightweight and installation-free, limiting each code package to 2MB for WeChat platform specifically. To extend their functionality, MiniApps use dynamic or on-demand loading of sub-packages (W3C, 2023a). During a cold-start or initialization, only the main package is loaded, with corresponding sub-packages loaded dynamically as users access specific pages.
Page Routing of MiniApps. WeChat MiniApps are multi-functional entities, with each page facilitating specific tasks and interconnected via page routing. Once initiated, WeChat client establishes a page stack, enabling page manipulations for the MiniApp. Two methods facilitate this routing process: the <navigator> widget in the render layer and the navigation-specific sub-app APIs in the logic layer. The former offers both intra-MiniApp and cross-MiniApp navigation, governed by the widget’s target attribute, with a variety of routing methods based on the open-type attribute. The latter implements routing in the logic layer, using the url property to specify the target page, with APIs like wx.navigateTo and wx.redirectTo enabling different navigational outcomes.
Event-driven Callbacks. MiniApps, being event-driven, respond to UI interactions with appropriate logic execution and interface updates. Consequently, static analysis of MiniApps should consider the UI event-driven callback functions that influence the control flow. This work focuses on two following crucial types of callbacks:
$\bullet$ Lifecycle Callbacks include two-pronged MiniApp lifecycle—App instance and Page instance lifecycles. For instance, a cold-started MiniApp first triggers the onLaunch callback of the App instance, and when the MiniApp launches and surfaces, it triggers the onShow callback. A page’s first load activates the onLoad callback of the Page instance, which in turn triggers the onShow callback when displayed and stacked.
$\bullet$ GUI Event Handler Callbacks facilitate communication between a MiniApp’s rendering and logic layers. Events, tied to UI components/widgets, trigger these callbacks when users interact, enabling GUI feedback processing.
2.2. Comparison with Native/Web Apps
In contemporary application development, MiniApps, Native Apps, and Web Apps represent three distinct paradigms. Although there are similarities among them, each one possesses its unique features and functionalities. In the following, we elucidate the differential features of these platforms, particularly focusing on two primary aspects:
Packaging and Building Mechanism: Unlike web apps that lack a specific packaging format, MiniApps and native apps require specific file formats for distribution and installation. MiniApps, commonly packaged as WXAPKG, diverge notably from native apps, which utilize APK or IPA formats for Android and iOS platforms, respectively. The building process of these applications also presents significant differences. Unlike native apps where the bundling process often allows access to all source code at once, the subpacking mechanism in MiniApps inherently limits static acquisition of code to only the main package. Subpackage code in MiniApps is not immediately available; it requires dynamic traversal and is loaded as needed, a stark difference from the comprehensive code availability seen in native app bundling.
Rendering Mechanism: MiniApps leverage a hybrid rendering approach specifically designed for WeChat ecosystem, combining WebView and Native components. Its unique components, such as <navigator>, akin to a hybrid of HTML’s <a> tag and native app navigation functions, exemplifies integration with WeChat’s user experience. Media components like <image>, <video>, and <camera> in WXML, while functionally similar to their counterparts in Native and Web Apps, are fine-tuned for performance and direct integration with WeChat platform.
Table 1. Comparison of MiniApps, Native Apps, and Web Apps.
Mechanism MiniApps Native Apps Web Apps
Packaging WXAPKG APK (Android) /
IPA (iOS)
Building Subpacking Bundling Bundling
Rendering Hybrid Rendering Native Rendering HTML/CSS Rendering
(WebView & Native components)
Rendering Code WXML XML (Android) HTML
Storyboard & XIB (iOS)
Logic Code JavaScript Java (Android) JavaScript
Swift or Objective-C (iOS)
Distribution WeChat App Store, Google Play Website Access
  • 1
    Note that Mech. is short for Mechanism.
3. Motivating Example
We provide a running example showing how MiniApp collects and uses privacy in Figure 2, which consists of three key stages: (1) Cross-package page routing. The user triggers cross-package page routing via <navigator> (❶) and <botton> (❹❺) widgets in pages/myInfo (main package), respectively navigating to pages/takePhoto (main package) and pages/checkID (subpackage dynamically loaded); (2) Event-driven callback execution. The user triggers binding events in pages/takePhoto (❷) and pages/checkID (❻) through UI interactions, and the logic layer (JavaScript) executes the event-driven callback functions takePhoto and onInput respectively; (3) Sensitive data flow propagation. In the asynchronous function callbacks, the sensitive data flow further completes the propagation (❸❼). Notably, in this example, due to the lack of sufficient page routing and binding events, the pages/demo in the sub-package and the startRecord function within it are considered to be unreachable, which could potentially leads to FPs in previous DFA-only approaches (Wang et al., 2023b; Meng et al., 2023; Li et al., 2023).
Figure 2. Code snippet of cross-file/cross-package privacy practices. Subfigure (a) and (b) illustrate cross-file privacy practices in the initially loaded main package, subfigure (a) and (c) represent cross-subpackage privacy practices (i.e., FNs in previous static-only approaches), and subfigure (d) demonstrates dead code without entry points (i.e., FPs generated in previous DFA-only approaches).
3.1. Challenges for Hybrid Analysis of MiniApps
Considering the unique nature of MiniApps, which relies on Web technologies and integrates with the capabilities of native apps, the hybrid analysis of MiniApps primarily faces two key challenges:
Challenge#1: Complexity of cross-file/package callback control flow resolution. In MiniApps, a prominent challenge emerges from the complexity in analyzing control flow, particularly due to the widespread practice of reusing callback functions across different files and subpackages. This aspect of MiniApp development present a unique scenario, which is not directly analogous to existing techniques in Android/iOS environments. Developers often define callback function logic in shared JavaScript files and subsequently import these callbacks into various pages and subpackages of the MiniApp. This results in a complex, intertwined control flow graph of cross-file and cross-(sub)package callbacks, significantly complicating the process of tracing user interactions and understanding event-driven behaviors.
Challenge#2: Inadequate in locating WeChat-specific components in the UI widgets tree The distinctive components within WXML introduce unique challenges in the realm of UI exploration. These challenges primarily stem from the limitations of traditional testing tools. Designed specifically to align with the WeChat ecosystem, WXML components diverge significantly from the typical Document Object Model (DOM) structures or native UI frameworks familiar to conventional tools. This divergence renders tools like UIAutomator ineffective in identifying and locating these components directly from the GUI widgets tree. This limitation necessitates the development or adaptation of specialized testing tools and methodologies that can cater to the unique structure and behavior of WXML components, ensuring effective and efficient testing of MiniApps.
3.2. Insights for Potential Solutions
Upon evaluating the complexities inherent in MiniApps, particularly considering the aforementioned challenges, we distill two refined insights for potential solutions.
Insight#1: Constructing fine-grained and comprehensive topological structures for MiniApps. To address Challenge#1, the primary focus is on constructing a detailed and comprehensive topological structure of MiniApps. This structure would facilitate a more nuanced analysis across files and subpackages, supporting an in-depth control (❶❷) and data flow (❸) analysis. Notably, integrating callback control flow analysis within this framework is crucial for dynamic UI exploration. It could also potentially be a starting point for static data-flow analysis, thereby reducing false positives in static analysis. This insight underscores the need for a multifaceted approach, blending fine-grained structural mapping with advanced flow analysis, to provide a robust foundation for page transition analysis and automated UI exploration.
Insight#2: Enhancing existing dynamic testing tools with fuzzy matching for platform-specific components In response to Challenge#2, the strategy involves augmenting existing dynamic testing tools to better handle platform-specific MiniApp widgets. This enhancement would be achieved by implementing a fuzzy matching mechanism based on key component attributes like name, type, and text. By establishing an accurate correlation between static WXML component and dynamic UI Widget Tree elements, we can significantly improve the identification of MiniApp-specific widgets during dynamic runtime. This approach not only ensures the correct recognition of unique MiniApp components but also enhances the overall effectiveness of dynamic testing tools. This insight highlights the importance of adapting testing methodologies to the specificities of the MiniApp platform, leveraging both static and dynamic analysis attributes for comprehensive testing coverage.
4. Methodology
In response to the two challenges, we propose MiniScope, an automated UI exploration and privacy inconsistency analysis framework based on two-phase iterative hybrid program analysis, as shown in Figure 3. The framework consists of three key modules: MiniApp Dependency Graph Generator (§4.1), Directed UI Explorer (§4.2), and Privacy Inconsistency Detector (§4.3), which further includes a policy analyzer and a privacy practice monitor, thereby facilitating flow-to-policy cross-validation. Our methodology unfolds in two distinct phases, each involving a round of static analysis followed by a round of UI exploration. The entire workflow is structured as follows:
Main Package Analysis (Phase One). The initial phase of MiniScope is dedicated to dynamically loading subpackages through UI exploration to obtain the full package. Firstly, MiniScope initializes the MiniApp Dependency Graph (MDG) for the main package. Subsequently, based on the guidance of MDG, the Directed UI Explorer employs a subpackage-directed breadth-first traversal strategy to explore subpackage pages, thus obtaining the full package of the MiniApp.
Full Package Analysis (Phase Two). In the second phase, MiniScope tackles the precise identification of privacy practices based on the analysis of merged full package. Mirroring the approach of the previous phase, we commence with constructing the complete MDG through static analysis. During the UI exploration stage, MiniScope adopts a privacy-practice-directed depth-first strategy, which is instrumental in conducting a runtime exploration of sensitive behaviors, allowing for a nuanced understanding of the MiniApp’s privacy practices.
Figure 3. The architecture and workflow of MiniScope.
4.1. MiniApp Dependency Graph Generator
To facilitate automated UI exploration and more precise privacy practice detection, MiniScope analyzes the UI state transition, callback control flow and universal data flow. By combining them, MiniScope is able to derive a thorough and accurate topological view of a MiniApp.
4.1.1. Analyzing UI State Transitions
A MiniApp comprises numerous pages, each performing diverse functions. Interactions and switching between these modules are realized through page routing. To precisely capture the page routing information, MiniScope constructs a UI State Transition Graph (UTG) to model this page routing logic.
Definition 1 (UTG): UI State Transition Graph (UTG) represents the static GUI model (Yang et al., 2018; Dong et al., 2018; Liu et al., 2022), which depicts page transition sequences in a MiniApp. Formally, the UTG can be expressed as a directed graph $G_{U}=(V_{U},E_{U},\lambda_{U})$, where $V_{U}$ is a set of GUI states of pages with properties (including type, resource ID, text, bounds, GUI event, and callback method); $E_{U}$ is a set of GUI events triggering page routing that act as the transition condition edges, and $\lambda_{U}$ is the function for parsing the page routing conditions.
UTG construction. During UTG construction, MiniScope analyzes the rendering and logic layers separately: (1) For the rendering layer, MiniScope focuses on the <navigator> widgets in the WXML file responsible for page navigation, with the open-type attribute signifying the routing method. (2) For the logic layer, MiniScope chiefly evaluates JavaScript code related to page routing APIs. For instance, wx.navigateTo adds the current page to the page stack and navigates to a new page. By examining five methods in the <navigator> widget and five types of page routing APIs in the logic layer as shown in Table 2, MiniScope can delineate all possible page transitions in the parsing function $\lambda_{U}$ and statically construct the UI state transition graph $G_{U}$.
Figure 4. A simplified UTG example of Express MiniApp. Corresponding to Figure 2, in Event 1, the user triggers ontap event of <navigator>, leading to a transition from pages/myInfo to pages/takePhoto; in Event 2, the user triggers ontap event of <button text="manual input">, and the corresponding event handler callback function naveToCheckID is executed, naveToCheckID callback function, resulting in a transition from pages/myInfo to pages/checkID.
A UTG example. Figure 4 illustrates a simplified UTG example of the Express MiniApp. In this example, the set $V_{U}$ contains three pages: “pages/myInfo/index”(A), “pages/takePhoto/index”(B), and texttt“pages/checkID/index”(C). In the MiniApp framework, all page widgets must be registered in the WXML file. We determine page states $V_{U}$ by traversing through these widgets. For example, PageA consists of six <navigator> widgets, and when we click on the <navigator url=“/pages/takePhoto/index”>, it will navigate to the corresponding page, making the widget a transition condition of the set $E_{U}$ from PageA to PageB.
Table 2. MiniApp page routing methods considered in UI state transition resolution.
Category Property/Method Description
¡navigator¿ navigate Keep the current page and navigate to a non-tabBar one.
navigateBack Close the current page and go back.
redirect Close the current page and navigate to a non-tabBar one.
reLaunch Close all pages and open a new one.
switchTab Close all pages and open a target tabBar page.
API wx.navigateTo Keep the current page and navigate to a non-tabBar one.
wx.navigateBack Close the current page and go back to a previous one.
wx.redirectTo Close the current page and navigate to a non-tabBar one.
wx.reLaunch Close all pages and open a new one.
wx.switchTab Close all pages and open a target tabBar page.
4.1.2. Analyzing Callback Control Flows
In MiniApps, rendering layer widgets bind to specific UI events. Upon occurrence of these events, the logic layer responds by executing the corresponding event handler callback function. Hence, we build the Callback Control Flow Graph (CCFG) to conduct callback function level control flow analysis.
Definition 2 (CCFG): Callback Control Flow Graph (CCFG) models sequences of GUI event callbacks or lifecycle callbacks (Yang et al., 2015). Formally, CCFG can be expressed as a directed graph $G_{C}=(V_{C},E_{C},\lambda_{C})$, where $V_{C}$ is a set of functions in the page of MiniApps, $E_{C}$ is a set of GUI events or lifecycle events that trigger callback functions that act as trigger condition edges, and $\lambda_{C}$ is the function for parsing the triggering conditions of callback functions.
Determining CCFG entry point. Prior to constructing, it is necessary to determine the CCFG entry points which are event-driven callback functions. This is accomplished by performing cross-language analysis on rendering layer widgets and their associated event handler callback functions. User interactions with these widgets trigger the corresponding event handling logic, activating the associated control flows within the MiniApp. As an example shown in Figure 2 (❷), a widget with “bindtap = takePhoto” attribute will bind to the Tap event, and execute the takePhoto function when a Tap event is triggered. Furthermore, the page lifecycle callbacks can also serve as CCFG entry points as they are tied to specific lifecycle events, executing relevant logic when these events occur.
CCFG construction. From the identified entry points, MiniScope constructs the rest of the function call chain within the JavaScript language to obtain a complete CCFG. Considering the highly customizable, framework-based nature of MiniApps, a tailored cross-file/package callback control flow analysis approach is necessary. Function definitions within MiniApp pages resemble class method definitions, as they’re all defined within the Page() object instance. During initialization, developers provide all lifecycle callback functions, event handler functions, and custom functions to the Page object as attribute methods. Method intercommunication requires the “this” keyword, referring to the current Page object. Notably, some commonly used callback functions can be defined in shared JavaScript files and then imported across various pages. Leveraging these observations, we design a cross-file callback control flow analysis algorithm based on the JavaScript Abstract Syntax Tree (AST). This algorithm takes into account the unique patterns of function usage and invocation in MiniApps, including the shared use of callback functions across different files and subpackages.
Definition 3 (AST): Abstract Syntax Tree (AST) (Yamaguchi et al., 2014) can represent the language structure of source code in detail. Formally, the AST can be expressed as an undirected property graph $G_{A}=(V_{A},E_{A},\lambda_{A},\mu_{A})$, where $V_{A}$ is a set of AST nodes, $E_{A}$ is a set of AST edges labeled by the function $\lambda_{A}$. Additionally, we assign a property to each node using $\mu_{A}$, which represents whether the node is an operator or an operand.
Construction algorithm. The main steps of cross-file CCFG construction are described in Algorithm 1. The algorithm takes inputs including the AST $G_{A}$ of the current page, a set of $pageMethods$, a set of $entryPoints$, and a set of $sensiAPIs$. The output of the algorithm is the complete CCFG $G_{C}$. In particular, the algorithm consists of the following three key components:
$\bullet$ Function getImportedMethods(): Iterates through all nodes in the AST. If a node is an $ImportDeclaration$, the $specifier$ of the node is appended to the $pageMethods$ list (line 3-5).
$\bullet$ Function getCCFG(): Iterates through the children of $func$. If a child node is a $CallExpression$, it retrieves the callee of the node (line 11-12). If the callee is a $MemberExpression$ and its left child is a $ThisExpression$ and its right child is in $pageMethods$, it adds an edge to $G_{C}$ and recursively calls $getCCFG$ for this node (line 13-18). If the callee is in $sensiAPIs$, it also adds an edge to $G_{C}$ (line 20-22).
$\bullet$ Main Process: Initializes $pageMethods$ by calling $getImportedMethods$ with the AST $G_{A}$ and an initial list of $pageMethods$ (line 26). Then iterates through the children of the AST $G_{A}$. If a child is a $CallExpression$, it gets the callee of the node (line 28-29). If this node is in $entryPoints$, it calls $getCCFG$ with the node, $pageMethods$, and an initial CCFG $G_{C}$ (line 30-32). Finally, it returns the constructed CCFG $G_{C}$.
Input: AST $G_{A}$ of the page, a set of $pageMethods$, a set of $entryPoints$, a set of $sensiAPIs$
Output: CCFG $G_{C}$ of the page
1 Function getImportedMethods($G_{A},pageMethods$):
2       for $node\in G_{A}.nodes$ do
3             if $node$ is $ImportDeclaration$ then
4                   $pageMethods.append(node.specifier)$
5             end if
6
7       end for
8      return $pageMethods$
9 end
10Function getCCFG($func$, $pageMethods$, $G_{C}$):
11       for $node\in func.children$ do
12             if $node$ is $CallExpression$ then
13                   $callee\leftarrow node.callee$
14                   if $callee$ is $MemberExpression$ then
15                         if $callee.leftchild$ is $ThisExpression$ and $callee.rightchild\in pageMethods$ then
16                               $G_{C}.addEdge(func,node)$
17                               getCCFG($node$, $pageMethods$, $G_{C}$)
18
19                         end if
20
21                   end if
22
23             end if
24            if $callee\in sensiAPIs$ then
25                   $G_{C}.addEdge(func,node)$
26
27             end if
28
29       end for
30      return $G_{C}$
31 end
[…]
33 for $node\in G_{A}.children$ do
34       if $node$ is $CallExpression$ then
35             $callee\leftarrow node.callee$
36             if $node\in entryPoints$ then
37                   getCCFG($node$, $pageMethods$, $G_{C}$)
38             end if
39
40       end if
41
42 end for
return $G_{C}$
Algorithm 1 Cross-file CCFG Construction
4.1.3. Merging into MinApp Dependency Graph
To precisely model MiniApps, MiniScope combines UTG and CCFG with UDFG proposed in TaintMini (Wang et al., 2023b) to create a unified topological structure of MiniApps, called MiniApp Dependency Graph (MDG). For example, MiniScope can analyze sensitive APIs distribution and their triggering paths by querying the graph. In particular, we first formally describe UDFG following its original definition, then we define MDG accordingly.
Definition 4 (UDFG): Universal Data Flow Graph (UDFG) can be used to capture the cross-language and cross-page data flow within a MiniApp (Wang et al., 2023b). Formally, UDFG can be represented as a directed graph $G_{D}=(V_{D},E_{D},\lambda_{D})$, where $V_{D}$ is the set of data objects, $E_{D}$ is the set of edges representing sensitive data flow dependencies and propagation directions. Correspondingly, $\lambda_{D}$ is the function for tracing data flow propagation.
Definition 5 (MDG): MiniApp Dependency Graph (MDG) is a joint topological structure combined by the above four directed graphs. The key insight for constructing MDG is that in each page’s CCFG and UDFG, a node exists for each statement and predicate in the source code. It is natural to merge the page node of UTG with the corresponding root node of each page’s AST. Therefore, MDG is formally represented as $G_{M}=(V_{M},E_{M},\lambda_{M})$, where:
  • $V_{M}=\sum V_{A}\cup V_{U}$
  • $E_{M}=\sum E_{A}\cup E_{U}\cup\sum E_{C}\cup\sum E_{D}$
  • $\lambda_{M}=\lambda_{A}\cup\lambda_{U}\cup\lambda_{C}\cup\lambda_{D}$.
MDG of the code snippet in the motivating example. As shown in Figure 5, to analyze the access to camera resources, MiniScope first locates the page “pages/takePhoto” (❶ Green Line) which is navigated through the <navigator> widget in “pages/index”. Then MiniScope determines the entry point of the CCFG is “takePhoto”, which is triggered by the Tap event bound to the <button> in the WXML (❷ Purple Line). Through the combination analysis of control flow and data flow, MiniScope can determine that the takePhoto function further calls wx.createCameraContext to create the context, which is then assigned to the variable “ctx”. Later, the camera is triggered to take a photo through the “ctx.takePhoto” method (❸ Pink Line).
Figure 5. MiniApp Dependency Graph (MDG) representation of the code snippet presented in Figure 2. The Black lines represent AST edges; the Green Lines represent UTG edges; the Purple Lines represent CCFG edges; and the Pink Lines represent UDFG edges.
4.2. Directed UI Explorer
Our dynamic analysis component leverages the constructed MDG to guide task-directed UI exploration. In the two-phase analysis, we respectively follow the subpackage-directed breadth-first exploration strategy and the privacy-practice-directed depth-first exploration strategy, thus achieving a balance between maximizing the UI state space and minimizing exploration time.
4.2.1. Fuzzing Matching between WXML Component and UI Widget Tree
As described in Challenge#2, due to the platform-specific widgets defined in WXML (e.g., <navigator>) and the support for programmable attributes (e.g., loop rendering with wx:for), WXML components lack the unique resourceID and deterministic XPATH in the rendered UI screen, making it difficult to locate them uniquely. Therefore, to leverage the static MDG for guiding dynamic UI exploration, we propose a fuzzy matching approach to establish a robust mapping between WXML component and UI widget tree. The fuzzy matching comprises two primary strategies, as outlined below, to accurately locate widgets on the rendered UI screen.
Key Attribute Matching. The first strategy involves the identification of a set of key attributes for each platform-specific WXML component, such as name, type, and text, etc. Then MiniScope employs the Intersection over Union (IoU) metric to match these key attributes with those of the widgets on the UI screen. The IoU is calculated as $IoU=\frac{W\cap U}{W\cup U}$, where $W$ represents a set of key attributes in the WXML component and $U$ represents a set of key attributes in the current UI secreen widgets.
XPATH Matching. In scenarios where UI widgets do not possess sufficient attribute information, or when the maximum IoU value between all widgets on the UI screen and the target WXML component falls below a 50% threshold (Yang et al., 2022a), MiniScope resorts to calculating the Levenshtein distance between the XPATH of target widget in WXML and widgets on the UI screen. The widget with the shortest Levenshtein distance is then selected as the candidate, followed by a verification of the UI state of the target page.
Through this fuzzy matching approach, MiniScope effectively bridges the gap between static MDG and dynamic UI exploration. By leveraging key attributes and the Levenshtein distance, MiniScope is able to accurately map WXML components to their corresponding widgets in the UI Widget Tree, overcoming the inherent challenges in locating platform-specific widgets within WXML.
4.2.2. Task-directed UI Exploration Strategy
MiniScope employs distinct exploration strategies in the two phases, meticulously designed to comprehensively analyze MiniApps by dynamically loading sub-packages and investigating privacy practices.
Subpackage-directed BFS Exploration. In Phase One, to effectively load sub-packages and maximize the exploration of page state space, MiniScope utilizes the MDG of the main package for BFS exploration. This process begins with extracting all sub-package page transition paths from the MDG and maintaining them in a queue. Leveraging Appium (appium, 2023), MiniScope obtains the current UI screen’s widgets and calculates the IoU and Levenshtein distance to accurately locate the target widget within the transition path. Once navigation is executed, MiniScope verifies if the current page path aligns with the expected one. Upon reaching a target sub-package page, all related page transition paths are removed from the queue, streamlining the exploration process.
Privacy-practice-directed DFS Exploration. Once BFS is completed, MiniScope unpacks and merge the dynamically loaded sub-package into the main package. A subsequent round of static analysis updates the MDG. In Phase Two, guided by the complete MDG, MiniScope delves into a DFS exploration focusing on privacy-related practices within the MiniApp. Similar to BFS, this exploration process involves computing GUI events on specific pages to trigger relevant callbacks. However, in contrast to BFS, the DFS exploration prioritizes completing each potential privacy practice path in a first-in, first-out order from the queue, ensuring a thorough examination of privacy practices.
Through the aforementioned two-phase task-directed UI exploration strategy, MiniScope achieves a comprehensive and systematic analysis of MiniApps, ensuring a robust and detailed understanding of MiniApp privacy practices, addressing the challenges posed by their unique features.
4.3. Privacy Inconsistency Detector
MiniScope could enable many security and software engineering tasks seamlessly. In this paper, our primary focus centers on the application of MiniScope to detect privacy inconsistency within MiniApps. Thus we employ an innovative approach to monitor the privacy practices within MiniApps by instrumenting sensitive APIs and cross validate with the statement in privacy policies to detect inconsistency.
4.3.1. Privacy Practice Monitor
In this component, MiniScope leverages the Frida framework (frida, 2023) to hook sensitive APIs. This enables us to monitor the API calls made by the MiniApp during dynamic UI exploration. When these API calls are triggered, our monitor captures crucial context information and stack traces, logging any detected statements. However, MiniApps execute their logic layer primarily through APIs that are encapsulated from native library functions via the JSBridge mechanism (Lu et al., 2020). Thus the main challenge we faced lies in the mapping between each sub-app API and the corresponding system API in Java layer. By reverse engineering WeChat, we found that invoking a MiniApp API from JavaScript triggers a native bridge function offered by the host platform, which manages the execution of the actual system APIs. In the context of WeChat MiniApps, this bridge function is located within the commonjni.AppBrandJSBridgeBinding class, specifically in the invokeCallbackHandlerI(int, java.lang.String) method. Utilizing objection (sensepost, 2023), a runtime mobile exploration toolkit powered by Frida (frida, 2023), MiniScope hooks this bridge function, checking its parameters and runtime call stack. This allows us to establish the mapping from each sub-app API to the corresponding Java layer system API (detailed in our online documentation (Anonymous, 2023)). Note that this process of constructing the WeChat-to-Android API Mapping is offline and can be reused throughout our analysis once built.
4.3.2. LLM-based Privacy Policy Analysis
To determine the disclosure of personal data collection and usage, we have drawn from the previous works of Polisis (Harkous et al., 2018) and PolicyLint (Andow et al., 2019), then construct three ontologies to describe the privacy statement.
Definition of Privacy Statement. As shown in Table 3, we represent privacy policies as a series of triples, i.e., (DC, SSoC, DE). Data Controller is the party responsible for determining the purposes and means of personal data processing, which can be the application itself (first party) or a third party. SSoC Verbs refer to a list of verbs that describe the Storing-Sharing-or-Collection of data. Data Entity represents any privacy information that can identify or reflect the activities of natural persons.
Table 3. Four types of ontologies in privacy policy.
Labels Ontology
DC Data Controller (First Party or Third Party)
SSoC Storing-Sharing-or-Collection Verbs
DE Data Entity (Privacy Information or Sensitive System Resources)
Prompt Design. Typically, previous works (Harkous et al., 2018; Andow et al., 2019; Ling et al., 2022) rely on manually annotated corpus to train named entity recognition models (such as CRF, BiLSTM, BERT, etc.) for predicting entity labels, or by constructing data and entity dependency (DED) trees to extract tuple representations of privacy policy statements. With the recent emergence of large language models (LLMs), we aim to accomplish privacy policy extraction through fine-tuning LLMs using few-shot in-context learning. The prompt consists of three parts: (1) Task Description, (2) Few-shot Demonstrations, and (3) Query. For better understanding, we provide the following prompt example.
Task Description: Here is a table where each row has 3 items. The first item is ”privacy detail”, and the second to third items are ”subject” and ”verb” for the ”privacy detail” in the same row. Here are some sample rows in the table: Few-shot Demonstrations: ”To save photos, the developer will request your permission to access your photo album”, ”developer”, ”access”, ”photo album” Query: Please read the table and follow the above sample rows to complete the table by filling ”subject”, ”verb”, ”privacy detail” of each row: {Input Sentence}
4.3.3. Privacy Inconsistency Analysis
API-to-Privacy Mapping. To bridge the semantic gap between API calls and privacy policies, especially concerning data entities, we collect detailed descriptions of 29 sensitive APIs from the official WeChat documentation (Wec, 2023c), including API functions, parameters, etc. We correlate the API usage to specific privacy data access/collection based on semantics. That is, the data entities in privacy policies are divided into 13 categories, as detailed in our online documentation (Anonymous, 2023).
Flow-to-Policy Cross-validation. If sensitive data flows like collection, storage, or sharing are explicitly disclosed in the privacy policy, we consider them policy-compliant. Inconsistencies arise otherwise. Drawing from the consistency model in PoliCheck (Andow et al., 2020), we categorize inconsistent privacy disclosures into two: 1) Redundant Disclosures (RD), where the policy mentions more than actual privacy-related practices. This could mean developers disclose all possible sensitive behaviors, mitigating potential risks. 2) Omitted Disclosures (OD), where privacy-related practices occur but are not mentioned in the policy.
5. Evaluation
We have implemented a prototype of MiniScope using 9,466 lines of code (LoC) with Python, excluding any third-party libraries or open-source tools. We then conducte an evaluation on MiniScope, aiming to answer the following research questions (RQs):
$\bullet$ RQ1 (Effectiveness): How effective is MiniScope in detecting privacy-related practices and analyzing privacy policies?
$\bullet$ RQ2 (Ablation Study): How does each component of MiniScope affect the performance separately?
$\bullet$ RQ3 (A Large-scale Study in the Real World): How prevalent are privacy compliance violations in MiniApps ecosystem?
5.1. Experimental Setup
Dataset. To collect MiniApps, we utilize the open-source tool MiniCrawler (Zhang et al., 2021) to download MiniApp packages from the WeChat App Market. We have collected a total of 127,460 MiniApps, with 289 GB of total size. To collect privacy policies from our list of MiniApps, we designe and deploy a privacy policy crawler. It requests privacy policies from the WeChat server based on the AppID, completing the collection process within 8 hours. Due to the prevalent absence of privacy policies, we find that only 10,786 (8.4%) MiniApps have valid privacy policies.
Ground Truth. To ensure a reliable evaluation of RQ1 and RQ2, we form a benchmark by randomly sampling 150 MiniApps from the dataset. They are dynamically executed for sub-package loading and manually inspected by three experienced researchers separately, which forms the ground truth of their privacy-related practices. To ensure a thorough dynamic interaction, each researcher interacts with the MiniApp for 5 minutes or until all reachable pages/functions are fully triggered, whichever happens first.
Running Environment. Our experiments are conducted on a server running Ubuntu Linux of 22.04 version with two 64-core AMD EPYC 7713 and 256 GB RAM. The dynamic testing is performed on an Android Virtual Device (AVD) with a system version of Android 8.1.0 and API version Level 27. The version of WeChat used is 8.0.37, and the WebView kernel version is 107.0.5304.141.
5.2. Effectiveness of MiniScope (RQ1)
The efficacy of MiniScope, as outlined in Section 4, is reliant on two key elements: 1) the accuracy of hybrid analysis; and 2) the effectiveness of privacy policy analysis. Our assessments of MiniScope focus on these two aspects.
Effectiveness of Hybrid Analysis. We compare the performance of MiniScope with that of TaintMini (Wang et al., 2023b) on our ground truth dataset. For a comprehensive assessment, we use the following metrics: True Positives (TPs), False Positives (FPs), False Negatives (FNs), precision, recall, and F1 score.
The results of the evaluation are listed in Table 4. MiniScope significantly surpasses TaintMini in detecting 19 out of 20 sensitive APIs, offering the highest precision, recall and F1 score. The lower precision of TaintMini stems from false positives found in unreachable code, while its decreased recall is due to its omission of sub-packages and its use of a single data flow analysis. This results in TaintMini overlooking many device-specific API calls like wx.openBluetoothAdapter, which may be called only once, not involving data flow propagation. MiniScope, by incorporating page state transition analysis, callback control flow analysis, and universal data flow analysis, provides the best overall performance.
Table 4. Effectiveness of hybrid analysis using MiniScope. Prerepresents Precision%; Recrepresents Recall%; F1% represents F1-score%.
Category APIs GroundTruth TaintMini MiniScope
TP FP FN PreRecF1% TP FP FN PreRecF1%
Location wx.chooseLocation 61 60 16 1 78.9 98.4 87.6 61 0 0 100.0 100.0 100.0
wx.getLocation 168 131 9 37 93.6 78.0 85.1 167 0 1 100.0 99.4 99.7
wx.onLocationChange 22 12 0 10 100.0 54.5 70.6 17 1 4 94.4 81.0 87.2
wx.startLocationUpdateBackground 9 1 0 8 100.0 11.1 20.0 4 1 4 80.0 50.0 61.5
Media wx.chooseImage 202 178 21 24 89.4 88.1 88.8 202 1 0 99.5 100.0 99.8
wx.chooseMedia 13 13 2 0 86.7 100.0 92.9 13 3 0 81.3 100.0 89.7
wx.chooseMessageFile 22 20 1 2 95.2 90.9 93.0 21 0 1 100.0 95.5 97.7
wx.chooseVideo 23 16 0 7 100.0 69.6 82.1 20 0 3 100.0 87.0 93.0
OpenAPI wx.chooseAddress 86 76 2 10 97.4 88.4 92.7 84 0 2 100.0 97.7 98.8
wx.chooseInvoiceTitle 7 5 0 2 100.0 71.4 83.3 6 0 1 100.0 85.7 92.3
wx.getUserInfo 60 38 8 22 82.6 63.3 71.7 60 0 0 100.0 100.0 100.0
wx.getUserProfile 171 146 2 25 98.6 85.4 91.5 168 0 3 100.0 98.2 99.1
wx.getWeRunData 22 13 0 9 100.0 59.1 74.3 15 2 5 88.2 75.0 81.1
Device wx.addPhoneContact 15 2 0 13 100.0 13.3 23.5 14 0 1 100.0 93.3 96.6
wx.createCameraContext 17 6 0 11 100.0 35.3 52.2 15 0 2 100.0 88.2 93.8
wx.createLivePusherContext 4 0 0 4 0.0 0.0 0.0 3 0 1 100.0 75.0 85.7
wx.getRecordManager 6 0 0 6 0.0 0.0 / 6 0 0 100.0 100.0 100.0
wx.openBluetoothAdapter 22 1 0 21 100.0 4.5 8.7 10 0 12 100.0 45.5 62.5
Album wx.saveImageToPhotosAlbum 175 85 16 90 84.2 48.6 61.6 168 1 7 99.4 96.0 97.7
wx.saveVideoToPhotoAlbum 14 4 0 10 100.0 28.6 44.4 12 0 2 100.0 85.7 92.3
Total / 1119 807 77 312 91.3 72.1 80.6 1066 9 49 99.2 95.6 97.4
Effectiveness of LLM-based Privacy Policy Analysis. We compare the LLM-based approach with five baseline NER models. Particularly, we utilize GPT-3.5-Turbo to perform inquiries. During this process, we keep the default configuration of GPT-3.5-Turbo, with $temperature=1$ and $top\_n=1$. To complete the experiment, we have utilized an estimation of 62 thousand tokens in total to analyze 100 privacy policies. For the other five baseline models, we train them using annotations from the CA4P-483 (Zhao et al., 2022) corpus. To ensure unbiased evaluation, we apply 10-fold cross-validation to the models, with 8 folds used for training, 1 fold for parameter tuning and optimization, and 1 fold for testing. Table 5 presents the detailed performance of these models. We observe that GPT-3.5-Turbo outperforms other baseline models significantly, indicating its high robustness and effectiveness.
Table 5. Comparison with baseline NER models. DC represents Data Entity; SSoC represents Storing-Sharing-or-Collection Verbs; DE represents Data Entity. Prerepresents Precision%; Recrepresents Recall%; F1% represents F1-score%.
Model DC SSoC DE
PreRecF1% PreRecF1% PreRecF1%
HMM 21.4 73.9 33.2 27.9 82.9 40.4 43.0 74.6 54.6
CRF 64.2 52.9 58.0 72.9 72.0 72.4 74.2 77.1 75.6
BiLSTM 61.0 65.3 63.1 78.6 79.4 78.9 80.4 75.0 77.6
BiLSTM-CRF 66.9 68.7 67.8 79.0 80.0 79.4 82.0 76.8 79.3
BERT 65.5 68.1 66.3 77.1 85.6 81.1 79.3 82.1 80.7
GPT-3.5-Turbo 100 100 100 95.4 91.4 93.4 100 91.8 95.7
For hybrid analysis, MiniScope outperforms RegMatch and TaintMini on our ground truth dataset with 99.2% precision, 95.6% recall, and 97.4% F1 score. Similarly, in privacy policy analysis, our LLM-based method surpasses existing NER models with 98.5% precision, 94.5% recall, and 96.5% F1 score.
5.3. Ablation Study (RQ2)
We conduct an ablation study to understand how UTG and CCFG improve MiniScope’s performance. Three MiniScope variants are developed for this: 1) MiniScope-UDFG-ONLY, which disables UTG and CCFG to emulate the TaintMini approach; 2) MiniScope-STATIC-ONLY, which enables UTG and CCFG; and 3) MiniScope-DYNAMIC-ONLY, which removes static guidance.
Number of Privacy-related Practice Detected by MiniScope Variants Figure 6 illustrates the number of privacy-related practices detected by MiniScope and three variant baselines on the ground truth dataset. It can be observed that MiniScope is a superset of each variant baseline. Furthermore, by combining static and dynamic approaches, MiniScope identifies 13 privacy-related practices that cannot be detected by each individual component. This indicates that the methodology adopted by MiniScope improves the performance of privacy-related practice identification and also demonstrates its robustness.
Figure 6. Venn diagram showing the privacy-related practices detected by each component of MiniScope.
Normalized Privacy-related Practice Detection by MiniScope Variants For better visualization and understanding, we presents the normalized privacy-related practices detected by MiniScope and other three variant baselines in Figure 7. The normalized value is computed as follows: $nomalized\_privacy\_practice=\frac{pp}{pp(\textsc{MiniScope})}$.
From the comparison between UDFG-ONLY and STATIC-ONLY, we notice a significant performance boost in static privacy-related practice identification due to the integration of UTG and CCFG, especially in the Device and Album categories. This reinforces our RQ1 findings where TaintMini’s dependence on single data flow analysis lead to increased false negatives in these API categories. STATIC-ONLY generally outperforms DYNAMIC-ONLY. This is a result influenced by MiniApps’ characteristics, such as the need for user login or specific purchase completions, which impact dynamic analysis recall. By utilizing hybrid analysis, MiniScope outperforms STATIC-ONLY and DYNAMIC-ONLY, efficiently combining their strengths for more effective privacy practice identification.
By conducting an ablation study on each component, we discover that the hybrid analysis employed by MiniScope indeed enhances the performance of privacy-related practice identification while also demonstrating robustness.
Figure 7. The performance of MiniScope, MiniScope-UDFG-ONLY, MiniScope-STATIC-ONLY and MiniScope-DYNAMIC-ONLY on normalized privacy-related practice identified.
5.4. A Large-scale Study in the Real World (RQ3)
Measurement. We utilize MiniScope to conduct a large-scale compliance violation detection on 127,460 MiniApps in the real world. The preliminary results indicate that out of these MiniApps, 33,570 (26.37%) collect and use privacy information. Howerer, only 10,786 (8.46%) of these MiniApps have valid privacy policies in place. This highlights the severe privacy risks prevalent in the MiniApp ecosystem. Furthermore, we assess the consistency of privacy-related practices against privacy policies to detect privacy compliance violations.
Table 6 presents the detailed statistical results of the compliance detection. Redundant Disclosure refers to disclosures in a privacy policy that exceed privacy-related practices, while Omitted Disclosure refers to disclosures in a privacy policy that are fewer than privacy-related practices. The former is considered a weak violation, while the latter is considered a strong violation. Due to the possibility of multiple instances of RD or OD in a MiniApp, we classify and record the distribution of disclosures for each type of privacy-related practice. Overall, we summarize our findings as follows.
Table 6. Privacy inconsistencies detected by MiniScope.
Category Scopes Total Redundant Omitted
Count PercentCount Percent%
Location UserLocation 7,714 1,389 18.1 64 0.9
Media ChooseMedia/File 13,115 1,594 12.2 193 1.5
OpenAPI Address 8,714 1,063 12.2 126 1.5
Invoice 3,540 490 13.9 7 0.2
UserInfo 14,535 1,380 9.5 116 0.8
WeRun 2,297 489 21.3 8 0.4
Device PhoneContact 1,179 163 13.9 24 2.1
PhoneCalendar 796 395 49.7 / /
Camera 8,283 809 9.8 9 0.2
Record 4,773 849 17.8 1 0.1
Bluetooth 1,941 494 25.5 13 0.7
Clipboard 54 15 27.8 / /
Album PhotoAlbum 11,198 1,095 9.8 250 2.3
Total / 78,139 10,225 13.1 811 1.1
Findings. In our study of compliance violations in MiniApps, we find that Redundant Disclosure (13.1% of all cases) significantly outweighs Omitted Disclosure (1.1%). MiniApps tend to excessively disclose privacy-related practices linked to PhoneCalendar (49.7%) more often, possibly due to developers retaining policy disclosures even after removing associated privacy practices once time-limited events conclude. In contrast, Omitted Disclosure is more prevalent in categories like Media, Address, UserInfo, and PhotoAlbum, likely due to developers not realizing these categories require permission as per WeChat’s policy (Wec, 2023c). Our large-scale study reveals that 5.7% (614/10786) of MiniApps over-collect private data secretly, while 33.4% (3599/10786) overstate their actual data collection.
Responsible Disclosure to Developers. Based on our findings, we have responsibly disclosed these violations to their developers via the email obtained from the privacy policies. Particularly, we shared our methodology and the trigger paths of potential privacy compliance violations detected by MiniScope with developers, and asked for their feedback on these findings. Overall, we notified a total of 2,282 developers, out of which 1,727 emails are successfully sent, 396 emails are intercepted by the server’s filtering policies, and 159 of them are reported as undeliverable email addresses.
Table 7. Responses from developers.
Response Measures Taken Count by Each Type Total
Acknowledgement accept and update the privacy policy overclaim (25) & over-collection (17) 42
partially accept and update the privacy policy overclaim (2) 2
Disagreement claim their reasons overclaim (2) 2
As shown in Table 7, as of the time of writing, we have received 46 responses from MiniApp developers. We summarize their responses as follows:
  • 42 developers fully accept our findings and commit to updating their privacy policies. Most acknowledge their privacy policies are outdated, with redundant disclosures stemming from past versions that incorporated relevant privacy practices (which further confirms our previous findings). Developers express a desire for the MiniApp platform to offer automated code audits and maintain consistent privacy compliance.
  • 2 developers partially accept our findings and provide valuable suggestions. They attribute some redundant disclosures to user-input data collection via forms in MiniApps. Thus, they consider our findings partially accurate, intending to selectively update based on our results. These insights underscore potential areas of improvement in our approach, particularly in dealing with user-input privacy within MiniApps.
  • 2 developers disagree with our findings. These developers highlight that a portion of the detected privacy over-claims in our study stemm from selective functionality accessibility, which is specifically tailored for certain target users or for time-limited promotions. This selective accessibility can be implemented based on special entry points, such as QR code scanning, leading to the scenario that our analysis tool is unable to access restricted pages within the MiniApps, thereby omitting the analysis of corresponding privacy practices. We further discuss such kind of detection corner cases in Section 6.1.
Case Studies. We conduct case studies inspecting detected privacy violation MiniApps to understand the concrete impacts of these violations. Here, we present two typical cases.
  • Redundant Disclosure. An example of extreme redundant disclosure is an online ordering MiniApp named WanMoTang(AppID: wx5e4dc66b2f***), with a 4.2-star rating and over 1k recent users. Although its privacy policy discloses nearly all privacy types (19 in total), the MiniApp only employs 3 types in practice: UserInfo, Location, and Photoalbum. This over-disclosure potentially escalates privacy risks, as any privacy practice may seem justifiable against an overly comprehensive privacy policy.
  • Omitted Disclosure. One instance of omitted disclosure in the MiniApp HaiHuiShou(AppID: wxded0379a82***), which has an average 4.0 rating and around 200 reviews. Although it discloses 4 privacy types in its policy, our analysis uncovered undisclosed practices in its pages/demo/demo source code. The onLoad function creates a Camera context, requests WeRun data, and collects Address information. These undisclosed practices constitute privacy compliance violations.
6. Discussion
6.1. Threats to Validity
Despite the significant contributions, our work still nonetheless carries some limitations that worth noting.
Influence from unpacking of MiniApps. Despite our attempts to utilize state-of-the-art open-source tools such as wxappUnpacker (xdmjun, 2023) and unveilr (r3x5ur, 2023), 235 (0.2%) of the MiniApps in our dataset cannot be successfully unpacked. Certain MiniApps, developed using third-party multi-end frameworks for a uniform programming style, may not have a standard page structure after unpacking. This situation, affecting around 1.3% of our ground truth dataset, contributes to an approximately 11.0% false negative rate in static-only analysis.
Corner cases encountered in the dynamic analysis. The effectiveness of the dynamic analysis could be hampered by the unexpected corner cases. In cases like online shopping MiniApps, where we cannot actually complete orders or access review pages, we depend on static-only analysis to detect potential privacy inconsistency. Meanwhile, as some developers have reported, some certain pages or functions are dynamically made available to users based on specific criteria or provided through special entry points (such as QR code scanning), which makes our static analysis unable to determine callback entries, and the dynamic analysis, accordingly, unable to perform.
Possible bias in ground truth. The ground truth used in our evaluation (RQ1 and RQ2) may be subject to manual confirmation bias. Despite our efforts to involve three experienced security researchers to conduct meticulous inspections and reach a consensus, there may still be potential biases, which could arise in sub-package dynamic loading or privacy-related practices identification.
6.2. Scalability and Transferability
In our research, while the primary focus is on the WeChat MiniApp ecosystem similar to other works (Wang et al., 2023b; Meng et al., 2023; Li et al., 2023), it is important to emphasize the scalability and transferability of our proposed methodology to other platforms. As highlighted in the W3C MiniApp Standardization White Paper (W3C, 2023a), MiniApps across most platforms share similar underlying architectures, employing JavaScript for logic programming and analogous layout files (WXML in WeChat, AXML in Alipay, SWAN in Baidu, and TTML in TikTok). This architectural uniformity suggests that MiniScope, initially tailored for WeChat, holds significant potential for broader applicability with minimal adjustments. By fine-tuning it to accommodate the nuances of each platform’s specific components and APIs, MiniScope can be effortlessly migrated and applied to other ecosystems beyond WeChat. This transferability not only enhances the utility and reach of our research but also opens avenues for comprehensive privacy and security analysis across diverse MiniApp platforms.
7. Related Work
MiniApps, as a novel application paradigm, have begun to attract scholarly attention in recent times. Prior studies in this field can be categorized into three primary aspects.
Security and Privacy of MiniApps. Several investigations have delved into the security aspects of Miniapps (Lu et al., 2020; Yang et al., 2022b; Zhang et al., 2022; Wang et al., 2022, 2023c; Zhang et al., 2023c; Liu et al., 2020). Notably, existing study (Wang et al., 2022) gathered 83 MiniApp bugs from real-world scenarios and developed WeDetector to identify WeBugs following three bug patterns. Another project (Lu et al., 2020) probed into issues like system resource exposure, subwindow deception, and subapp lifecycle hijacking within the Mini-Program ecosystem. They conducted evaluations on 11 popular platforms to ascertain the widespread nature of these security problems. Besides, a novel privacy leak issue in MiniApps is studied (Zhang et al., 2022) that could potentially lead to private data theft by the MiniApp platform. They elucidated an attack process exploiting this vulnerability. Lastly, a Cross Mini-Program request forgery vulnerability (CMRF) (Yang et al., 2022b) is figured out in Mini-Program communication. They shed light on the root cause and designed the CMRFScanner tool for its detection. Additionally, a series of studies have emphasized the importance of privacy in MiniApp ecosystem (Yang et al., 2023; Zhang et al., 2023b; Baskaran et al., 2023; Wang et al., 2023b; Meng et al., 2023; Zhang et al., 2023a; Wang et al., 2023a; Li et al., 2023; Wang et al., 2023d). TaintMini (Wang et al., 2023b) introduced a framework for detecting flows of sensitive data within and across mini-programs using static taint analysis. Another work MiniTracker (Li et al., 2023) constructed assignment flow graphs as common representation across different host apps and performed a large-scale study on 150k MiniApps, which revealed the common privacy leakage patterns. Moreover, several studies (Meng et al., 2023; Baskaran et al., 2023; Zhang et al., 2023b) have focused on taint analysis technique to detect AppSecret leaks. In particular, another work (Wang et al., 2023a) focused on the consistency of data collection and usage in MiniApps. They crawled 2,998 MiniApps and detected 89.4% of them violated their privacy policies. More recently, Zhang et al. introduced SPOChecker and performed the first systematic study of privacy over-collection in MiniApps. As listed in Table 8, despite these significant contributions to understanding privacy dimensions, these static-only and DFA-only approaches fall short in accurate privacy behavior identification, indicating a crucial area for further exploration and research.
Table 8. Comparison of AutoMini with other MiniApp analysis tools. Note that ○ indicates that the tool does not support this feature; ◑ signifies that the tool supports this feature, but there is a gap when compared to the SOTA techniques; ● denotes the SOTA techniques available.
Approach Static Analysis Dynamic Analysis
UI Transition CFA DFA & Taint Subpackage UI Exploration Instrumentation
TaintMini (Wang et al., 2023b)
MiniTracker (Li et al., 2023)
WeMinT (Meng et al., 2023)
SPOChecker (Zhang et al., 2023a)
Wang et al. (Wang et al., 2023a)
WeJalangi (Liu et al., 2020)
AutoMini
Privacy Analysis of Various Platforms. The exploration of privacy analysis in various platforms, particularly the alignment between actual app behaviors and stated privacy policies, has become increasingly pivotal in recent years. Considerable efforts have been directed towards assessing privacy policy adherence in mobile apps, which is often performed either automatically by analyzing sensitive API calls (Slavin et al., 2016; Zimmeck et al., 2016) or user inputs (Wang et al., 2018; Nan et al., 2015). Tools like PoliCheck (Andow et al., 2020) have considering the recipients of personal data, thereby improving the accuracy of compliance analysis. Additionally, PurPliance (Bui et al., 2021) detects the data-usage purposes inconsistencies between privacy policy and the actual behavior of Android apps. Beyond mobile applications, researchers have delved into other platforms. For instance, ExtPrivA (Bui et al., 2023) and Ling et al. (Ling et al., 2022) focus on detecting inconsistencies between the privacy disclosures and data-collection behavior in browser extensions. OVRseen (Trimananda et al., 2022) compares network traffic and privacy policies to analyze personal data exposed by OVR apps. While previous works have provided valuable insights into privacy compliance across various platforms, our research extends these efforts to the emerging field of MiniApps.
8. Conclusion
Our paper presents MiniScope, a tool for detecting privacy inconsistency using hybrid analysis. MiniScope leverages static analysis for high-level guidance in dynamic testing and cross-validates runtime behavior with privacy policies. Evaluations show MiniScope identifies privacy practices with an average precision, recall, and F1 score of 99.2%, 95.6%, and 97.4% respectively. Our large-scale study reveals privacy inconsistency issues in MiniApp ecosystem, with only 10,786 out of 127,460 WeChat MiniApps providing valid privacy policies, and 2,282 (21.2%) displaying various privacy violations. Our findings have been responsibly disclosed to 2,282 developers, with 44 acknowledgments. We believe MiniScope will aid researchers and developers in identifying and mitigating privacy risks in MiniApps.